what value of x satisfies the system of equations below
x+2y=11
4x-y=8

Answer by ewatrrr(24785)


Hi
x+2y=11
4x-y=8     0r y = 4x-8

x+2y=11
8x-2y=16
    9x = 27
     x = 3    and  y = 4      +y+=+12-8 
P(3,4) is the solution for this system of equations

Wish You the Best in your Studies.